### v1

@@ 6.4.1 Methodology
Complementing the overall Chamfer distance, we employ a specific evaluation protocol to assess the reconstruction accuracy on fine geometric details, which can be averaged out by global metrics. Therefore, as previously we focus not only on global reconstruction, but also on high curvature areas. However, the DTU objects being less “curvy” than other benchmarks, high curvature areas were segmented slightly differently, in order to focus on fine structures.
@@ 6.4.2 Results
However, our results are much stabler in the sparse-view scenario, as indicated in Table 11 where the number of views is decreased from 49 to 5. Our results are also much more accurate in high curvature areas, as indicated in these tables. Notably, although we use Colmap’s normals, our CDs are $10\%$ lower in these areas, which highlights the interest of the proposed normal integration framework. Overall, these MVS experiments provide a proof of concept for the incorporation of our framework beyond MVPS.
@@ 7 Conclusion and Perspectives
We presented a versatile and efficient framework for high-fidelity 3D surface reconstruction by leveraging multi-view normal cues, with optional use of reflectance information. Through a novel reparametrization of normals and reflectance as radiance vectors simulated under varying illumination, our method fits seamlessly into both traditional MVS and NVR pipelines. It is also very fast, thanks to CUDA acceleration and an optimizable hash grid. Extensive experiments on the most recent benchmarks demonstrate state-of-the-art performance, particularly in capturing fine details and handling challenging reflectance and visibility conditions.
However, challenges remain. We observe bias in the volume rendering version when using ground truth normals, and robustness to noisy input normals still needs improvement. These limitations underline the importance of better PS estimation, especially with a limited number of light sources.
@@
Robin Bruneau’s postdoctoral fellowship was funded by the Department of Quantitative Biomedicine at the University of Zurich. Baptiste Brument’s doctoral student fellowship was funded by the CNRS through the project Open-DOPAMIn. Lilian Calvet’s postdoctoral fellowship was supported by the University of Zurich, the University Hospital of Balgrist, and the OR-X - a Swiss national research infrastructure for translational surgery. The research leading to these results received funding from the French National Research Agency (ANR) through the LabCom project ALICIA-Vision, and by the Department of Computer Science at the University of Copenhagen (DIKU) through the Copenhagen Data+ project PHYLORAMA.

### v3

@@ 6.4.1 Methodology
Complementing the overall Chamfer distance, we employed a specific evaluation protocol to assess the reconstruction accuracy on fine geometric details, which can be averaged out by global metrics. Therefore, as previously we focus not only on global reconstruction, but also on high curvature areas. However, the DTU objects being less “curvy” than other benchmarks, high curvature areas were segmented slightly differently, in order to focus on fine structures.
@@ 6.4.2 Results
However, our results are much stabler in the sparse-view scenario, as indicated in Table 11, where the number of views is decreased from 49 to 5. Our results are also much more accurate in high curvature areas, as indicated in these tables. Notably, although we use Colmap’s normals, our CD are $10\%$ lower in these areas, which highlights the interest of the proposed normal integration framework. Overall, these MVS experiments provide a proof of concept for the incorporation of our framework beyond MVPS.
@@ 7 Conclusion and Perspectives
We presented a versatile and efficient framework for high-fidelity 3D surface reconstruction by leveraging multi-view normal cues, with optional use of reflectance information. Through a novel re-parametrisation of normals and reflectance as radiance vectors simulated under varying illumination, our method fits seamlessly into both traditional MVS and NVR pipelines. It is also very fast, thanks to CUDA acceleration and an optimisable hash grid. Extensive experiments on the most recent benchmarks demonstrate state-of-the-art performance, particularly in capturing fine details and handling challenging reflectance and visibility conditions.
However, challenges remain. We observe bias in the volume rendering version when using ground truth normals, and robustness to noisy input normals still needs improvement.
@@
Robin Bruneau’s postdoctoral fellowship was funded by the Department of Quantitative Biomedicine at the University of Zurich. Baptiste Brument’s doctoral student fellowship was funded by the CNRS through the OPEN-DOPAMIn project. Lilian Calvet’s postdoctoral fellowship was supported by the University of Zurich, the University Hospital of Balgrist, and the OR-X - a Swiss national research infrastructure for translational surgery. The research leading to these results received funding from the French National Research Agency (ANR) through the LabCom project ALICIA-Vision, and by the Department of Computer Science at the University of Copenhagen (DIKU) through the Copenhagen Data+ project PHYLORAMA.